40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 16:56:38 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades in progress 
    Reply  Wed Oct 9 17:27:35 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades  
       Reply  Wed Oct 9 22:05:55 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Fri Oct 11 17:37:08 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
             Reply  Sun Oct 13 13:29:09 2013, rana, Update, LSC, New lockin / sensing matrix screens 
                Reply  Mon Oct 14 17:40:15 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
                   Reply  Thu Oct 17 13:40:55 2013, Jenne, Update, LSC, New lockin / sensing matrix frequencies All_REFL_PRMI_17Oct2013.pdf
             Reply  Mon Oct 14 21:12:35 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Wed Oct 16 03:04:37 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
             Reply  Wed Oct 23 19:28:22 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
Message ID: 9250     Entry time: Thu Oct 17 13:40:55 2013     In reply to: 9237
Author: Jenne 
Type: Update 
Category: LSC 
Subject: New lockin / sensing matrix frequencies 

I locked PRMI, and (after fixing the POP QPD situation, noted in elog 9249) took power spectra of all the REFL RFPDs.  It looks like the area above 500 Hz is pretty clean and flat for all the signals, so I'm going to use 560Hz, 562Hz, 564Hz, 566Hz and 568Hz for my 5 sensing matrix frequencies.

Also, I'm not sure what is going on with REFL11, but there's a weird dip between 630 Hz and 660 Hz in both I and Q.  I recentered this guy not too long ago (elog 9218), but it clearly needs some more looking-at.

All_REFL_PRMI_17Oct2013.pdf

ELOG V3.1.3-